OpenStudy (anonymous):

How do I extract m(t) from this graph in order to apply it to this equation?

OpenStudy (anonymous):

\[\lim_{T \rightarrow \infty} \frac {1} {2T} \int\limits_{-T}^{T} m^{2} (t) dt\]

OpenStudy (anonymous):

Do I just find an equation for each line and then square it to have m(t) ^2 ?

OpenStudy (atlas):

yeah you are right.........there will be two equations for m(t)......one will be for 0<t<T/2 and other for T/2<t<T

OpenStudy (anonymous):

Thanks!

OpenStudy (atlas):

you are welcome............remember that whenever there is a discontinuous function - we always have to break equations like this.